Grade 11General Physics

The direction of 3 forces 20N, 40N, 60N acting at a point are parallel to sides of equilateral triangle taken in order . The magnitude of resultant vector isA 10√3NB 30NC 0 ND 20√3N

Munesh Antil
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
Lets put an x-axis on the direction of 10N force and a perpendicular y- axis.
The x – coordinates  of the resultant = 20 – 40 cos 60-60cos60 = -30N
The y – coordinates of the resultant = 40sin60-60sin60
 = 20√3-30√3 = -10√3
Size of resultant =sqrt ((30)^2 +(10√3)^2)
                            = sqrt(900+300) = sqrt(1200)
                            = 10√12N.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)